# Explainability of Machine Learning Models under Missing Data

### v1

@@ 3.2 Imputation techniques
    XGBoost (Extreme Gradient Boosting) [47] is a powerful and efficient algorithm that belongs to the family of gradient boosting techniques. It builds an ensemble of decision trees, where each tree corrects errors made by the previous ones, enhancing predictive accuracy. XGBoost stands out for its speed and performance, employing advanced features like tree pruning, regularization, and parallel processing, which help in reducing overfitting and handling large-scale data. XGBoost can handle missing data directly.
@@
    Mean Imputation is a classical technique used to handle missing data in a dataset. When some data points are missing, mean imputation replaces these missing values with the mean (average) of the available data for that particular variable. This method is straightforward and easy to implement, ensuring that all cases remain in the analysis and preserving the sample size.
@@
    Multiple Imputation by Chained Equations (MICE) [36] is a technique for imputing missing data in datasets through regression. It considers the variability of missing values by crafting various imputed datasets. This technique entails sequentially modeling each inadequate variable based on other data variables, producing numerous imputations for each lacking value. The iteration ceases when the divergence among different entries falls below a set threshold. Through this approach, MICE preserves inter-variable connections and yields a stronger and more precise approximation of the missing values.
@@
    Conditional Distribution-based Imputation of Missing Values with Regularization (DIMV) [37], is an innovative algorithm designed to handle missing data. This method works by determining the conditional distribution of a feature with missing entries, leveraging information from fully observed features. In addition, the DIMV algorithm is robust to the assumption of multivariate normal and does not require any label information. Therefore, it can be used for both supervised and unsupervised learning. DIMV can also explain the contribution of each feature to the imputation of a feature with a missing value in a regression coefficient-like manner. Moreover, the technique is robust to multicollinearity due to $L_{2}$ norm regularization.
@@
    SOFT-IMPUTE [40] is an imputation technique based on matrix completion and low-rank approximation, particularly useful for handling missing data in large and sparse datasets. This method operates by iteratively replacing the missing entries with values that minimize the reconstruction error of the data matrix, using singular value decomposition (SVD). Furthermore, SOFT-IMPUTE alternates between imputing missing values and performing a soft-thresholding operation on the singular values of the data matrix. This process encourages the resulting matrix to be low-rank, effectively capturing the underlying structure of the data while filling in the missing entries.
@@ 4 Theoretical analysis
In this section, we consider a simple linear regression model to show the effects of missing data in training and test data on the mean absolute Shapley values of a mean imputation method.
Assume we have a training data $\mathcal{D}=(\mathbf{x}|\mathbf{y})$ which has $n$ samples, $\mathbf{x}$ is univariate input, and $\mathbf{y}$ is a target. The corresponding missing data of $\mathcal{D}$ is denoted by $\mathcal{D}^{*}=(\mathbf{x^{*}}|\mathbf{y})$. Here, we assume that only $\mathbf{x}$ contain missing values and denote $\text{Obs}(\mathbf{x})$ as the index set of observed samples in $\mathbf{x}^{*}$, i.e., $\text{Obs}(\mathbf{x}^{*})=\{i\in\mathbb{N}:x_{i}\text{ is observed}\}$. Moreover, we denote $\mathcal{D}^{\prime}=(\mathbf{x}^{\prime}|\mathbf{y})$ as imputed data by using mean imputation. It means that
@@
$y^{\prime}=\widehat{f}(x^{\prime})=\beta^{\prime}_{0}+\beta^{\prime}_{1}x^{\prime},$
@@
$\beta_{1}=\frac{\text{Cov}(\mathbf{x},\mathbf{y})}{\text{Var}(\mathbf{x})},$
$\beta^{\prime}_{1}=\frac{\text{Cov}(\mathbf{x^{\prime}},\mathbf{y})}{\text{Var}(\mathbf{x^{\prime}})}.$
We consider a test data $\mathbf{z}$ which has $m$ samples. It should be reminded that the test data $\mathbf{z}$ could contain missing values (denoted by $\mathbf{z}^{*}$) and the imputed version $\mathbf{z}^{\prime}$ is presented as
@@
Then, we want to illustrate a global feature importance of $\mathbf{x}$ and $\mathbf{x}^{\prime}$ on $\mathbf{z}$ and $\mathbf{z}^{\prime}$ respectively by using mean absolute Shapley values, i.e.,
$|\phi(\mathbf{z})|=\frac{1}{m}\left(\sum_{j=1}^{m}|\phi(z_{j})|\right),$
$|\phi^{\prime}(\mathbf{z}^{\prime})|=\frac{1}{m}\left(\sum_{j=1}^{m}|\phi^{\prime}(z^{\prime}_{j})|\right).$
where $\phi(\mathbf{z})$ and $\phi(\mathbf{z^{\prime}})$ are the global feature importance on $\mathbf{z}$ and $\mathbf{z^{\prime}}$ respectively; $\phi(z_{j})$ and $\phi(z^{\prime}_{j})$ are the Shapley values on the prediction of sample $z_{j}$ and $z^{\prime}_{j}$ respectively.
The following theorem shows that the global feature importance on $\mathbf{z^{\prime}}$ could be simplified.
Theorem 1
@@
Proof 1
In [9], the Corollary 1, the Shapley value on the prediction of sample $z^{\prime}_{j}$ for $j=1,2,\dots,m$ with the corresponding model on and $\mathcal{D}^{\prime}$ are defined as
$\phi^{\prime}(z^{\prime}_{j})=(z^{\prime}_{j}-\mathbb{E}[\mathbf{x^{\prime}}])\beta^{\prime}_{1}.$
Now, by using the mean imputation (2), if $z_{j}$ is missing then $z^{\prime}_{j}=\mathbb{E}[\mathbf{x}^{*}]$. Thus,
$\phi^{\prime}(z^{\prime}_{j})=(\mathbb{E}[\mathbf{x}^{*}]-\mathbb{E}[\mathbf{x^{\prime}}])\beta^{\prime}_{1}.$
@@
$\phi^{\prime}(z^{\prime}_{j})=0,\text{ for }j\notin\text{Obs}(\mathbf{z}^{*}).$
@@
$=\frac{1}{m}\left(\sum_{j\in\text{Obs}(\mathbf{z}^{*})}|\phi^{\prime}(z^{\prime}_{j})|\right).$
@@
Based on the theorem, we have the following analysis. First, when the missing rate of $\mathbf{z}^{*}$ increases, the number of samples that have non-zero contributions in $|\phi^{\prime}(\mathbf{z}^{\prime})|$ is decreased. We can find that $|\phi^{\prime}(\mathbf{z}^{\prime})|$ would be small and less than $|\phi(\mathbf{z})|$. However, one important thing is that we cannot conclude that $|\phi^{\prime}(\mathbf{z}^{\prime})|$ is always less than $|\phi(\mathbf{z})|$. The reason is that $|\phi^{\prime}(z^{\prime}_{j})|$ could be greater than $|\phi(z_{j})|$ for some $j\in\text{Obs}(\mathbf{z}^{*})$. To be clearly, expanding the formulas $\phi(z_{j})$ and $\phi^{\prime}(z^{\prime}_{j})$ gives
$|\phi(z_{j})|=\left|(z_{j}-\mathbb{E}[\mathbf{x}])\beta_{1}\right|=\left|z_{j}-\mathbb{E}[\mathbf{x}]\right|.\frac{|\text{Cov}(\mathbf{x},\mathbf{y})|}{\text{Var}(\mathbf{x})},$
@@
$|\phi^{\prime}(z^{\prime}_{j})|=\left|(z^{\prime}_{j}-\mathbb{E}[\mathbf{x^{\prime}}]\beta^{\prime}_{1}\right|=\left|z^{\prime}_{j}-\mathbb{E}[\mathbf{x^{\prime}}]\right|.\frac{|\text{Cov}(\mathbf{x^{\prime}},\mathbf{y})|}{\text{Var}(\mathbf{x^{\prime}})}.$
@@
Note that if $j\in\text{Obs}(\mathbf{z}^{*})$ then $z^{\prime}_{j}=z_{j}$ and if we suppose that $\mathbb{E}[\mathbf{x}]\approx\mathbb{E}[\mathbf{x}^{\prime}]$, then the difference between $\phi(z_{j})$ and $\phi^{\prime}(z^{\prime}_{j})$ comes from the remaining terms $\frac{|\text{Cov}(\mathbf{x},\mathbf{y})|}{\text{Var}(\mathbf{x})}$ and $\frac{|\text{Cov}(\mathbf{x^{\prime}},\mathbf{y})|}{\text{Var}(\mathbf{x^{\prime}})}$. Additionally, these terms depend on the missing rate in training data $\mathbf{x}$ and $\mathbf{x}^{\prime}$. In particular, because $\mathbf{x^{\prime}}$ is the imputed version of $\mathbf{x}$ with mean imputation then $\text{Var}(\mathbf{x^{\prime}})\leq\text{Var}(\mathbf{x})$. It means when the missing rate in training data increases, the term $\frac{|\text{Cov}(\mathbf{x^{\prime}},\mathbf{y})|}{\text{Var}(\mathbf{x^{\prime}})}$ could be greater than $\frac{|\text{Cov}(\mathbf{x},\mathbf{y})|}{\text{Var}(\mathbf{x})}$. Thus, $|\phi^{\prime}(z^{\prime}_{j})|$ could be greater than $|\phi(z_{j})|$ for some $j\in\text{Obs}(\mathbf{z}^{*})$. In summary, there is a different trend in global feature importance changes, this depends on the missing rates in training and test sets. A high missing rate in the test data can increase the global feature importance, but in the training data, it reduces that value.
@@
Second, the distribution of those Shapley values on $\mathbf{z}$ and $\mathbf{z^{\prime}}$ should have a zero mean. The evidence for that is
$\mathbb{E}[\phi(z_{j})]$ $=(\mathbb{E}[\mathbf{z}]-\mathbb{E}[\mathbf{x}])\beta_{1},$
$\mathbb{E}[\phi(z_{j})]$ $=(\mathbb{E}[\mathbf{z}^{*}]-\mathbb{E}[\mathbf{x}^{\prime}])\beta^{\prime}_{1}.$
@@
In practice, $\mathbf{x}$ and $\mathbf{z}$ are split from one data, thus it may have $\mathbb{E}[\mathbf{z}]\approx\mathbb{E}[\mathbf{x}]$ and $\mathbb{E}[\mathbf{z}^{*}]\approx\mathbb{E}[\mathbf{x}^{\prime}]$. Therefore, $\mathbb{E}[\phi(z_{j})]\approx\mathbb{E}[\phi(z_{j})]\approx 0$. One important thing is that this assumption may be true for other imputation methods. This is why we see the symmetry in beeswarm plots of original data and mean imputation.
@@ 5.1 Experiment settings
In this study, we compare the Shapley values of the original data before missing data simulation with the Shapley values of XGBoost [47], a method that can be trained directly on missing data, and a two-step approach (imputation followed by regression). For the second alternative, during the imputation process, we test five methods: mean imputation, Multiple Imputation by Chained Equations (MICE) [36], conditional Distribution-based Imputation of Missing Values with Regularization (DIMV) [37], missForest [29], and SOFT-IMPUTE [40]. These methods are implemented with default settings using the fancyimpute  [^2] https://github.com/iskandr/fancyimpute, scikit-learn [48], and DIMV Imputation packages  [^3] https://github.com/maianhpuco/DIMVImputation.
We used the following data sets sourced from the Machine Learning Database Repository at the University of California, Irvine [^4] http://archive.ics.uci.edu/ml. Detailed descriptions of the datasets are provided in Table 1.
Dataset #Features #Samples
California 9 20,640
@@
We simulate missing values by randomly generating the ratio of missing entries to total entries in the features, with missing rates ranging from 20% to 80% in 20% increments. The experiments are run on a Windows-based machine equipped with an AMD Ryzen 7 3700X 8-Core Processor running at 3.59 GHz, 16 GB of RAM, and a 64-bit operating system. Each experiment is repeated ten times, and then the results are averaged. The codes for the experiments are available at https://github.com/simulamet-host/SHAP.
5.2 Global feature importance plot
We present the Global Feature Importance plots across various missing rates for the California dataset in Figures  123 and 4, while the plots for the diabetes dataset are included in A. However, the key points in the following inference on the California datasets can also be seen in the diabetes dataset.

### v2

@@ 3.2 Imputation techniques
    XGBoost (Extreme Gradient Boosting) [52] is a powerful and efficient algorithm that belongs to the family of gradient boosting techniques. It builds an ensemble of decision trees, where each tree corrects errors made by the previous ones, enhancing predictive accuracy. XGBoost stands out for its speed and performance, employing advanced features like tree pruning, regularization, and parallel processing, which help in reducing overfitting and handling large-scale data. XGBoost can handle missing data directly both during training and inference.
@@
    Mean Imputation (MI) is a basic technique consisting of replacing the missing values with the mean (sample average) of the available data for that particular variable. This method is straightforward and easy to implement.
@@
    Multiple Imputation by Chained Equations (MICE) [34] is a technique for imputing missing data in datasets through regression. It considers the variability of missing values by crafting various imputed datasets. This technique entails sequentially modeling each incomplete variable based on other data variables, producing numerous imputations for each lacking value. The iteration ceases when the divergence among different entries falls below a set threshold. Through this approach, MICE preserves inter-variable connections and yields a stronger and more precise approximation of the missing values.
@@
    Conditional Distribution-based Imputation of Missing Values with Regularization (DIMV) [35], is an innovative algorithm designed to handle missing data. This method works by determining the conditional distribution of a feature with missing entries, leveraging information from fully observed features. Even though DIMV is based on the assumption that the data follows a multivariate normal distribution, it is robust in the sense that it still works well when this assumption is not met. DIMV can also explain the contribution of each feature to the imputation of a feature with a missing value in a regression coefficient-like manner. Moreover, the technique is robust to multicollinearity due to $L_{2}$ norm regularization.
@@
    SOFT-IMPUTE [40] is an imputation technique based on matrix completion and low-rank approximation, particularly useful for handling missing data in large and sparse datasets. This method operates by iteratively replacing the missing entries with values that minimize the reconstruction error of the data matrix, using singular value decomposition (SVD). Furthermore, SOFT-IMPUTE alternates between imputing missing values and performing a soft-thresholding operation on the singular values of the data matrix. This process promotes a low-rank structure in the resulting matrix, effectively capturing the underlying low-rank structure of the data while filling in the missing entries.
  • 7.
    Generative Adversarial Imputation Nets (GAIN) [41] is a deep learning method for handling missing data by leveraging the Generative Adversarial Networks (GAN) framework. In GAIN, the generator (G) is trained to impute missing values in a data vector by conditioning the observed values, producing a completed vector that closely resembles the true data distribution. The discriminator (D), on the other hand, is tasked with distinguishing between observed and imputed components within this completed vector.
@@ 4 Theoretical analysis
In this section, we consider a simple linear regression model to analyze the effects of missing data (a missing-completely-at-random mechanism) in training and test datasets on global feature importance, measured by mean absolute Shapley values.
Assume we have an original training dataset, denoted as $\mathcal{D}=(\mathbf{x},\mathbf{y})$, consisting of $n$ samples, where $\mathbf{x}$ is a univariate input and $\mathbf{y}$ is the target variable. The corresponding dataset with missing values from $\mathcal{D}$ is denoted as $\mathcal{D}^{*}=(\mathbf{x^{*}},\mathbf{y})$. Here, only $\mathbf{x}^{*}$ has missing values and we denote $\text{Obs}(\mathbf{x}^{*})$ as the index set of observed samples in $\mathbf{x}^{*}$, i.e., $\text{Obs}(\mathbf{x}^{*})=\{i\in\{1,2,\dots,n\}:x_{i}\text{ is observed}\}$. Moreover, we denote $\mathcal{D}^{\prime}=(\mathbf{x}^{\prime},\mathbf{y})$ as imputed data by using mean imputation. It means that
@@
$y^{\prime}=\widehat{f^{\prime}}(x^{\prime})=\beta^{\prime}_{0}+\beta^{\prime}_{1}x^{\prime},$
@@
$\beta_{1}=\frac{\text{Cov}(\mathbf{x},\mathbf{y})}{\text{Var}(\mathbf{x})},\beta_{0}=\mathbb{E}[\textbf{y}]-\beta_{1}\mathbb{E}[\textbf{x}],\beta^{\prime}_{1}=\frac{\text{Cov}(\mathbf{x^{\prime}},\mathbf{y})}{\text{Var}(\mathbf{x^{\prime}})},\text{ and }\beta^{\prime}_{0}=\mathbb{E}[\textbf{y}]-\beta^{\prime}_{1}\mathbb{E}[\textbf{x}^{\prime}].$
We consider a test dataset $\mathbf{z}$, consisting of $m$ samples. It should be recalled that the test data $\mathbf{z}$ could contain missing values (denoted by $\mathbf{z}^{*}$) and the imputed version $\mathbf{z}^{\prime}$ is presented as
@@
Then, we want to illustrate a global feature importance of $\mathbf{x}$ and $\mathbf{x}^{\prime}$ on $\mathbf{z}$ and $\mathbf{z}^{\prime}$ respectively by using the mean absolute Shapley values, i.e.,
$|\phi(\mathbf{z})|=\frac{1}{m}\left(\sum_{j=1}^{m}|\phi(z_{j})|\right),$ (4)
$|\phi^{\prime}(\mathbf{z}^{\prime})|=\frac{1}{m}\left(\sum_{j=1}^{m}|\phi^{\prime}(z^{\prime}_{j})|\right),$ (5)
where $|\phi(\mathbf{z})|$ and $|\phi^{\prime}(\mathbf{z^{\prime}})|$ represent the global feature importance on $\mathbf{z}$ and $\mathbf{z^{\prime}}$, respectively; $\phi(z_{j})$ and $\phi^{\prime}(z^{\prime}_{j})$ denote the Shapley values for the predictions of samples $z_{j}$ and $z^{\prime}_{j}$. The following theorem shows that the global feature importance on $\mathbf{z^{\prime}}$ can be simplified.
Theorem 1.
@@
Proof.
In [9], the Corollary 1, the Shapley value for the prediction on sample $z^{\prime}_{j}$ for $j=1,2,\dots,m$ with the corresponding model on $\mathcal{D}^{\prime}$ are defined as
$\phi^{\prime}(z^{\prime}_{j})=(z^{\prime}_{j}-\mathbb{E}[\mathbf{x^{\prime}}])\beta^{\prime}_{1}.$ (7)
By using the mean imputation (3), if $z_{j}$ is missing then $z^{\prime}_{j}=\mathbb{E}[\mathbf{x}^{*}]$. Thus,
$\phi^{\prime}(z^{\prime}_{j})=(\mathbb{E}[\mathbf{x}^{*}]-\mathbb{E}[\mathbf{x^{\prime}}])\beta^{\prime}_{1},\text{ for }j\notin\text{Obs}(\mathbf{z}^{*}).$ (8)
@@
$\phi^{\prime}(z^{\prime}_{j})=0,\text{ for }j\notin\text{Obs}(\mathbf{z}^{*}).$ (9)
@@
$=\frac{1}{m}\left(\sum_{j\in\text{Obs}(\mathbf{z}^{*})}|\phi^{\prime}(z^{\prime}_{j})|\right).$ (10)
@@
Based on the theorem, we have the following analysis. First, when the missing rate of $\mathbf{z}^{*}$ increases, the number of samples that have non-zero contributions in $|\phi^{\prime}(\mathbf{z}^{\prime})|$ is decreased. We can find that $|\phi^{\prime}(\mathbf{z}^{\prime})|$ would be small and less than $|\phi(\mathbf{z})|$. However, one important thing is that we cannot conclude that $|\phi^{\prime}(\mathbf{z}^{\prime})|$ is always less than $|\phi(\mathbf{z})|$. The reason is that $|\phi^{\prime}(z^{\prime}_{j})|$ could be greater than $|\phi(z_{j})|$ for some $j\in\text{Obs}(\mathbf{z}^{*})$. To state it clearly, expanding the formulas $\phi(z_{j})$ and $\phi^{\prime}(z^{\prime}_{j})$ gives
$|\phi(z_{j})|=\left|(z_{j}-\mathbb{E}[\mathbf{x}])\beta_{1}\right|=\left|z_{j}-\mathbb{E}[\mathbf{x}]\right|.\frac{|\text{Cov}(\mathbf{x},\mathbf{y})|}{\text{Var}(\mathbf{x})},$ (11)
@@
$|\phi^{\prime}(z^{\prime}_{j})|=\left|(z^{\prime}_{j}-\mathbb{E}[\mathbf{x^{\prime}}]\beta^{\prime}_{1}\right|=\left|z^{\prime}_{j}-\mathbb{E}[\mathbf{x^{\prime}}]\right|.\frac{|\text{Cov}(\mathbf{x^{\prime}},\mathbf{y})|}{\text{Var}(\mathbf{x^{\prime}})}.$ (12)
@@
Note that if $j\in\text{Obs}(\mathbf{z}^{*})$, then $z^{\prime}_{j}=z_{j}$. Given that the missing mechanism is missing-completely-at-random (MCAR), we have $\mathbb{E}[\mathbf{x}]=\mathbb{E}[\mathbf{x}^{\prime}]$, so the difference between $|\phi(z_{j})|$ and $|\phi^{\prime}(z^{\prime}_{j})|$ comes from the remaining terms $\frac{|\text{Cov}(\mathbf{x},\mathbf{y})|}{\text{Var}(\mathbf{x})}$ and $\frac{|\text{Cov}(\mathbf{x^{\prime}},\mathbf{y})|}{\text{Var}(\mathbf{x^{\prime}})}$. Additionally, these terms depend on the missing rate in training data $\mathbf{x}$ and $\mathbf{x}^{\prime}$. In particular, because $\mathbf{x^{\prime}}$ is the imputed version of $\mathbf{x}$ with mean imputation then $\text{Var}(\mathbf{x^{\prime}})\leq\text{Var}(\mathbf{x})$. It means when the missing rate in training data increases, the term $\frac{|\text{Cov}(\mathbf{x^{\prime}},\mathbf{y})|}{\text{Var}(\mathbf{x^{\prime}})}$ could be greater than $\frac{|\text{Cov}(\mathbf{x},\mathbf{y})|}{\text{Var}(\mathbf{x})}$. To illustrate this, consider that, when the $i$-th sample becomes missing, the impact on the covariance is
$\text{Cov}(\mathbf{x^{\prime}},\mathbf{y})-\text{Cov}(\mathbf{x},\mathbf{y})=\frac{-1}{N}(y_{i}-\mathbb{E}[\mathbf{y}])\cdot(x_{i}-\mathbb{E}[\bf x^{\prime}]),$ (13)
which can be positive or negative and can contribute to an increase or decrease in the absolute value of the covariance. The derivation of the expression above is provided in Appendix B.
Thus, $|\phi^{\prime}(z^{\prime}_{j})|$ could be greater than $|\phi(z_{j})|$ for some $j\in\text{Obs}(\mathbf{z}^{*})$. In summary, there is a different trend in global feature importance changes, this trend depends on the missing rates in training and test datasets. A high missing rate in the test dataset can decrease global feature importance, while a high missing rate in the training dataset can increase it.
@@
Second, the distribution of the Shapley values on $\mathbf{z}$ and $\mathbf{z^{\prime}}$ should have a zero mean. The evidence for that is
$\mathbb{E}[\phi(\mathbf{z})]$ $=\mathbb{E}[(\textbf{z}-\mathbb{E}[\mathbf{x}])\beta_{1}]=(\mathbb{E}[\mathbf{z}]-\mathbb{E}[\mathbf{x}])\beta_{1},$ (14)
$\mathbb{E}[\phi^{\prime}(\mathbf{z^{\prime}})]$ $=\mathbb{E}[(\textbf{z}^{\prime}-\mathbb{E}[\mathbf{x}])\beta^{\prime}_{1}]=(\mathbb{E}[\mathbf{z}^{\prime}]-\mathbb{E}[\mathbf{x}^{\prime}])\beta^{\prime}_{1}.$ (15)
@@
In practice, $\mathbf{x}$ and $\mathbf{z}$ come from one dataset, so it is possible to have $\mathbb{E}[\mathbf{z}]\approx\mathbb{E}[\mathbf{x}]$ and $\mathbb{E}[\mathbf{z}^{\prime}]\approx\mathbb{E}[\mathbf{x}^{\prime}]$. Therefore, $\mathbb{E}[\phi(\mathbf{z})]\approx\mathbb{E}[\phi^{\prime}(\mathbf{z^{\prime}})]\approx 0$. This is why we see the symmetry in beeswarm plots of the original data and mean imputation.
@@ 5.1 Experiment settings
In this study, we conduct experiments to study the effects of missing values on Shapley values on regression and classification using four data sets from the Machine Learning Database Repository at the University of California, Irvine [^2] http://archive.ics.uci.edu/ml. Detailed descriptions of the datasets are provided in Table 1. Specifically, three scenarios were considered in our experiments: (1) Linear Regression on Original data (denoted as LRO and considered the ground truth), (2) XGBoost without imputation, and (3) linear regression on imputed data. Here, in the third scenario, we tested six imputation methods: Mean Imputation (MI), Multiple Imputation by Chained Equations (MICE) [34], conditional Distribution-based Imputation of Missing Values with Regularization (DIMV) [35], missForest [29], SOFT-IMPUTE [40], and Generative Adversarial Imputation Nets (GAIN) [41]. These methods are implemented with default settings using the fancyimpute  [^3] https://github.com/iskandr/fancyimpute, scikit-learn [53], and DIMV Imputation packages  [^4] https://github.com/maianhpuco/DIMVImputation. For regression tasks, the prediction model used is XGBoost regressor, while for classification, it is XGBoost Classifier. The code for the experiment is made available on Github.
We use global feature importance and beeswarm plots to analyze the effects of missing rate and handling missing strategy on Shapley values. Furthermore, the Mean Square Error (MSE) can also be utilized for that evaluation. For regression tasks, the MSE is calculated between the true values and the predicted values in the test set labels. For classification tasks, the MSE is computed between the imputed values and the original values (before simulating missing data) in the test set inputs. Similarly, MSESHAP is defined as the MSE between the Shapley values derived from a model fitted on the imputed data and the Shapley values derived from a model fitted on the original data. In the resulting tables and figures, MI denotes Mean Imputation, and SOFT denotes SOFT-IMPUTE. Values highlighted in boldface indicate the best performance.
Tasks Datasets #Features #Samples
Regression California 9 20,640
@@
For the missing data, we simulated missing-completely-at-random (MCAR) data from the original dataset by randomly assigning missing values with missing rates ranging from 0.2 to 0.8. The missing rate $r$ is defined as the ratio of the number of missing values to the total number of values in the dataset.
@@
The experiments were run on a Windows-based machine equipped with an AMD Ryzen 7 3700X 8-Core Processor running at 3.59 GHz, 16 GB of RAM, and a 64-bit operating system. Each experiment was repeated ten times, and the results were averaged. The code for the experiments is available at https://github.com/simulamet-host/SHAP
5.2 Global feature importance plot analysis
In this section, we analyze the effects of missing rate and handling missing strategy on global feature importance. We focus on examining the results of the California dataset because the key insights in this case can also be seen in the Diabetes dataset. We present the Global Feature Importance plots across various missing rates for the California dataset in Figures 123 and 4, while the plots for the Diabetes dataset are included in A. In addition, it is important to note that we compare eight processing workflows: LRO (Linear Regression on Original data, considered the ground truth), XGBoost, MI, MICE, DIMV, missForest, SOFT-IMPUTE, and GAIN. In the case of XGBoost, no previous imputation is necessary as that technique can deal with missing data by default. For the last six imputation methods, we simplify the notation by using the name of the imputation method to represent the combination of that method followed by linear regression.